OpenStudy (anonymous):

The average speed of an express train is 15mi/hr faster than the average speed of a local train. It takes the local train 2 hours longer than the express train to cover 360 miles. What is the average speed of the express train?

OpenStudy (anonymous):

let x=average speed of express train l=average speed of local train and t=time. \[l+15=x\]\[l=x-15\]\[\left( t+2 \right)*l=360\]\[t*x=360\]\[t=360/x\]\[\left( \left( 360/x \right)+2 \right)*l=360\]\[\left( \left( 360\over x \right)+2 \right)*\left( x-15 \right)=360\] \[x=60,x=-45\] average speed of express train is 60mph